OpenStudy (anonymous):

What is the y-intercept of a line through (2,7) and perpendicular to the line y= -3/2x +6. Is 6 and (2/3) correct?

OpenStudy (anonymous):

When perpendicular the slope is the opposite reciprocal (in this case it will be 2/3) y=mx+b will be your formula 7=(2/3)(2)+b Solve for b! then your formula will be y=(2/3)x + [what you got for b]

OpenStudy (anonymous):

Is 6 (2/3) right?

OpenStudy (mathmale):

Ryan is correct in showing that the new line will have the slope 2/3. But I think it very important to know how he got that. Anyone care to explain this?

OpenStudy (anonymous):

He did the opposite reciprocal . I need to know how to find the y-intercept and I think I got it but don't know if I did it correctly.

OpenStudy (anonymous):

well first its (2/3)(2) = 4/3 7 = 4/3 + b 21/3 = 4/3 + b (I changed 7 to a fraction to make it easier) SUBTRACT 4/3 FROM EACH SIDE 17/3 = b 5 and 2/3 = b

OpenStudy (mathmale):

You're right on in regard to how he got the slope of the perpendicular line. So: this perpendicular line with slope 2/3 goes through the given point (2,7). Substitute this info intothe point-slope formula \[y-y _{0}=m(x-x _{0})\] and solve for y. that will give you your y-intercept, b.

OpenStudy (anonymous):

Ok I got 5 and 2/3 . That's right?

OpenStudy (anonymous):

yes :)

OpenStudy (anonymous):

Thank ya (:

OpenStudy (mathmale):

Sassy: I'd suggest you actually do the work, so you'd know where that y-intercept b came from.